jhonyy9 (jhonyy9):

how many will be x from this equation ? 1 2,51 k ----- = -2lg( -------- + ------- ) sqrtx Re(sqrtx) 3,71D

OpenStudy (anonymous):

Please use the equation editor - that is why it is there.

jhonyy9 (jhonyy9):

why ? not is understandably ?

jhonyy9 (jhonyy9):

\[1/\sqrt{x} = -2\lg(2,51/(\Re(\sqrt{x})) + k/(3,71D)\]

jhonyy9 (jhonyy9):

\[(-\sqrt{x})/2x = \lg((9,31D +kRe(\sqrt{x}))/(3,71DRe(\sqrt{x}))\]